OpenStudy (anonymous):

Plz, I need help. How can I find an equation of the line passing through the two points. Write the equation in standard form. (-6,2) and (6,-9) Thx!!

OpenStudy (agreene):

start by using point-slope form, and then solve for y=mx+b form \[y-y_1=m(x-x_1)\] y-2=(-9-2)/(6-(-6))(x-(-6)) y-2=(-12/12)(x+6) y-2=-(x+6) y-2=-x-6 y=-x-4

OpenStudy (phi):

Here is a link to the various forms you can have for the equation of a line http://en.wikipedia.org/wiki/Linear_equation#Forms_for_2D_linear_equations Notice that your question asks for standard form.